One number is 1/4 of another number. The sum of the two numbers is 10. what are the two numbers?
Found 2 solutions by rfer, Maths68:
Answer by rfer(16322)

x+1/4x=10
1 1/4x=10
I am going to use decimals from here.
x=10/1.25
x=8
0.25x=2

Answer by Maths68(1474)


Let x be the one number and y be the other number
x is 1/4 of y
x=1/4(y)
x=y/4
4x=y
y=4x..........(1)

The sum of the two numbers is 10
x+y=10...........(2)

Put the value of y from (1) to (2)
x+4x=10...........(2)
5x=10
Divide by 5 both sides of above equation
5x/5=10/5
x=2

Put the value of x in (1)
y=4x..........(1)
y=4(2)
y=8
So one number is 2 and another number is 8